2. Show HOW to use inverses to solve the equation 7 -4 logs (x-3) = 6 for x. Show ALL algebraic steps!
Answer #1

Soln briven that 7-410g, (2-3) = 6 - 410g, (x-3)= 7-6 - 410g, (x-3)=1 - log, (x-3) = // 4 puing = (x-3) = gry log x= a = (1-3